Please help me with this:
It asks to have the inequality solved and to state and graph the solution set???
x/x-3<=-8/x-6 HELP PLEASE!!!

Answer by stanbon(75887)

state and graph the solution set???
x/(x-3)<=-8/(x-6)
Find the values x CANNOT take.
x cannot be 3 and x cannot be 6.
----------------
Draw a number line; plot 3 and 6 on that number line.
You now have three intervals of points that may be solution points.
Pick a test point from each interval to determine if the selected
interval is part of the solution set.
--------------
Interval (-inf,3); select x=0; 0/(0-3)<=-8/(0-6) ; true so (-inf,3) is part of the solution.
-----------
Interval (3,6); select x= 5; 5/(5-3)<=-8/(5-6) ; true so (3,6) is part of the
solution.
------------
Interval (6,inf); select x= 7; 7/(7-3)<=-8/(7-6); false so no solution here
---------------
Final solution: (-inf,3)U(3,6)
===============
Cheers,
Stan H.